If Z-N(0,1), show Var(Z) = 2.
Answer #1

TOPIC:Normal distribution.

=) Given that, 2 ~NO,D. The pdf of z is → e ; - 00<zco. can be . The even-onden moments of a obtained as — & [22] . 2n Jez) dDO 2 - 19 (2) = 2 2n - 2 Ž dz. Ag (-2) = (-2) e 2 2n - 24/2 ,00 = g(21. 2ne dz . : g(z) is an even junction of z. o gez) dz =Putting n=1 in we gety =) E (2²) = 21, r (1 + 77. [ r(3/2) I = 22. (1/2). 7:567 **** (? Again, E (24) . ( Puting ne2 in 2,5/2